# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

### v1

@@ 2.1 Convex formulation
where $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})$ is the $p\times p$ diagonal matrix of $\boldsymbol{\mathbf{\Omega}}$. Our model is to incorporate covariates as in (3) to $(\boldsymbol{\mathbf{\theta}},\boldsymbol{\mathbf{\Theta}})$ instead of $(\boldsymbol{\mathbf{\mu}},\boldsymbol{\mathbf{\Omega}})$, namely
$\boldsymbol{\mathbf{\theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}.$ (7)
To see that this leads to jointly convex nodewise regression problems, fix $j\in[p]$ and consider the partial regression of component $X_{j}$ against the covariates $\bm{U}$ and the remaining components $\bm{X}_{-j}$ in the general setting (1). For a matrix $\boldsymbol{\mathbf{M}}$ and sets of indices $\mathcal{I}$ and $\mathcal{J}$, denote by $\boldsymbol{\mathbf{M}}_{\mathcal{I},\mathcal{J}}$ the sub-matrix of $\boldsymbol{\mathbf{M}}$ consisting of rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$. Letting $-j$ indicate all indices excluding $j$, we have the conditional distribution
$X_{j}-\mu_{j}\>|\>\bm{X}_{-j},\bm{U}\sim N\left(\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j}),\,\boldsymbol{\mathbf{\Sigma}}_{jj}-\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}\boldsymbol{\mathbf{\Sigma}}_{-j,j}\right),$ (8)
suppressing the dependence of $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\mu}}$ on the covariates $\boldsymbol{\mathbf{u}}$ for notational convenience. Defining the error terms $\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}})$ where $\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$, by the matrix block inversion formula we may write (8) as the linear model
@@
which is readily seen to be jointly convex in the parameters $\theta_{j}$ and $\boldsymbol{\mathbf{\Theta}}_{j,-j}$ as defined in (6). In light of the specification of covariate dependence (7) and with $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and $\beta_{jkh}=[\boldsymbol{\mathbf{B}}_{h}]_{jk}$, the above leads to the nodewise regression model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}X_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}X_{k}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}})$ (9)
which allows for simultaneous estimation of the parameters $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\beta_{jkh}$ in a jointly convex setting.
Implicit in (6) is that $\Theta_{jj}=-1$ for all $j\in[p]$. This is satisfied in (7) by letting $[\boldsymbol{\mathbf{B}}_{0}]_{jj}=-1$ and $[\boldsymbol{\mathbf{B}}_{h}]_{jj}=0$ for all $h\in[q]$ and $j\in[p]$. By doing so, we assume as in Zhang and Li [2022] that the residual variance $\sigma^{2}_{\varepsilon_{j}}$ is not covariate-dependent, allowing us to write $\Omega_{jj}=\Omega_{jj}(\boldsymbol{\mathbf{u}})$ for $j\in[p]$.
@@ 3 Estimation
Suppose we collect $n$ independent observations $\{(\boldsymbol{\mathbf{x}}^{(i)},\boldsymbol{\mathbf{u}}^{(i)})\}_{i=1}^{n}$ of responses and covariates that follow the joint distribution in (1) with (6) and (7). Let $\boldsymbol{\mathbf{U}}=\left[\boldsymbol{\mathbf{u}}^{(1)},\dotsc,\boldsymbol{\mathbf{u}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times q}$ be the matrix of covariates and $\boldsymbol{\mathbf{X}}=\left[\boldsymbol{\mathbf{x}}^{(1)},\dotsc,\boldsymbol{\mathbf{x}}^{(n)}\right]^{\top}$ the matrix of responses with $\boldsymbol{\mathbf{x}}_{j}\in\mathbb{R}^{n}$ denoting the $j$-th column of $\boldsymbol{\mathbf{X}}$ and $\boldsymbol{\mathbf{u}}_{j}\in\mathbb{R}^{n}$ the $j$-th column of $\boldsymbol{\mathbf{U}}$. For $j\in[p]$, let $\boldsymbol{\mathbf{W}}_{-j}$ be the $n\times(p-1)(q+1)$ matrix of interactions of the remaining responses with the covariates. Concretely,
$\begin{split}\boldsymbol{\mathbf{W}}_{-j,0}&=\left[\boldsymbol{\mathbf{x}}_{1},\dotsc,\boldsymbol{\mathbf{x}}_{j-1},\boldsymbol{\mathbf{x}}_{j+1},\dotsc,\boldsymbol{\mathbf{x}}_{p}\right]\in\mathbb{R}^{n\times(p-1)},\\ \boldsymbol{\mathbf{W}}_{-j,h}&=\left[\boldsymbol{\mathbf{x}}_{1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{j-1}\odot\boldsymbol{\mathbf{u}}_{h},\ \boldsymbol{\mathbf{x}}_{j+1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{p}\odot\boldsymbol{\mathbf{u}}_{h}\right]\;\text{for $h\in[q]$},\\ \boldsymbol{\mathbf{W}}_{-j}&=\left[\boldsymbol{\mathbf{W}}_{-j,0},\ \boldsymbol{\mathbf{W}}_{-j,1},\ \dotsc,\ \boldsymbol{\mathbf{W}}_{-j,q}\right]\in\mathbb{R}^{n\times(p-1)(q+1)},\end{split}$ (10)
where $\odot$ denotes the elementwise product of two vectors. By writing
$\boldsymbol{\mathbf{\beta}}_{j,h}=(\beta_{j1h},\dotsc,\beta_{j,j-1,h},\beta_{j,j+1,h},\dotsc,\beta_{jph})^{\top}\in\mathbb{R}^{p-1}\;\text{and}\;\boldsymbol{\mathbf{\beta}}_{j}=(\boldsymbol{\mathbf{\beta}}_{j,0},\boldsymbol{\mathbf{\beta}}_{j,1},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q})^{\top}\in\mathbb{R}^{(p-1)(q+1)},$
we can view $\boldsymbol{\mathbf{\beta}}_{j}$ as $q+1$ blocks of $(p-1)$-element vectors $\boldsymbol{\mathbf{\beta}}_{j,h}$ for $h=\{0,1,\dotsc q\}$ so that (9) may be written in the block form
@@
Figure 1 relates the blocks $\boldsymbol{\mathbf{\beta}}_{j,h}$ to $\boldsymbol{\mathbf{\Theta}}$ in (7).
The joint convexity of (11) in $(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j})$ allows for simultaneous estimation of the penalized problem
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j})$ (12)
where $g_{j}$ is an arbitrary convex penalty function. Moreover, by the relationship between $\boldsymbol{\mathbf{\beta}}_{j}$ and rows of $\boldsymbol{\mathbf{\Theta}}$, sparsity-inducing penalties $g_{j}$ in (12) will give us sparse estimates of $\boldsymbol{\mathbf{\Theta}}$. From (6), we see that $\boldsymbol{\mathbf{\Theta}}$ has the same sparsity pattern as $\boldsymbol{\mathbf{\Omega}}$. Hence we keep the conditional independence interpretation of the sparsity in $\boldsymbol{\mathbf{\Theta}}$ as in $\boldsymbol{\mathbf{\Omega}}$.
@@
3.1 Estimating covariate-adjusted graphs
In the nodewise regression approach to GGMs, a symmetrization step is needed to ensure that the estimate of $\boldsymbol{\mathbf{\Omega}}$ is symmetric because each regression is fit independently [Meinshausen and Bühlmann, 2006]. In our approach, the nodewise regressions target $\boldsymbol{\mathbf{\Theta}}$, which is not necessarily symmetric due to the different scaling factor for each row in (6). To ensure symmetry in the estimate of $\boldsymbol{\mathbf{\Omega}}$, we perform a similar symmetrization step after performing all nodewise regressions. First, specify an estimate $\hat{\sigma}^{2}_{\varepsilon_{j}}$ of the error variance in (11). Then for an estimate $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$, set $\tilde{\boldsymbol{\mathbf{\beta}}}_{j}=-\hat{\boldsymbol{\mathbf{\beta}}}_{j}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ and define for all $h=0,1,\dotsc,q$ the symmetric matrix
@@ 3.1 Estimating covariate-adjusted graphs
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\left[\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\right]+\tilde{\beta}_{kjh}\left[\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\right],$ (13)
where the expression $\left[P\right]$ is equal to $1$ if $P$ is true and 00 otherwise. The mean vector and precision matrix may then be estimated by (6) via
@@
The preceding steps are summarized in Algorithm 1 in Section S1.
Equation (13) is the “and-rule” to symmetrize the matrices $\tilde{\boldsymbol{\mathbf{B}}}_{h}$ for $h=0,1,\dotsc,q$; $[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero if both $\tilde{\beta}_{jkh}$ and $\tilde{\beta}_{kjh}$ are nonzero. A less conservative estimate would be given by the “or-rule”, namely
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}{\left[\lvert\tilde{\beta}_{jkh}\rvert\geq\lvert\tilde{\beta}_{kjh}\rvert\right]}+\tilde{\beta}_{kjh}{\left[\lvert\tilde{\beta}_{jkh}\rvert\leq\lvert\tilde{\beta}_{kjh}\rvert\right]},$
so that $[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero if either $\tilde{\beta}_{jkh}$ or $\tilde{\beta}_{kjh}$ is nonzero. We elect to use the more conservative rule and note that both approaches are asymptotically equivalent [Meinshausen and Bühlmann, 2006] and that the and-rule has been considered before in covariate-adjusted graphical models [Cai et al., 2013, Zhang and Li, 2022].
3.2 Sparsity structure
With the application to eQTL studies in mind, we will focus on a particular sparsity-inducing penalty $g_{j}$ in (15). We wish to identify elementwise sparsity within a group $h$, amounting to sparsity in the coefficient vector $\boldsymbol{\mathbf{\beta}}_{j,h}$. We may interpret sparse entries in $\boldsymbol{\mathbf{\beta}}_{j,h}$ to mean that the covariate $u_{h}$ affects the conditional independence between $X_{j}$ and some, but not all other responses $X_{k}$. Meanwhile, we wish to identify groupwise sparsity, where $\boldsymbol{\mathbf{\beta}}_{j,h}=\boldsymbol{\mathbf{0}}$ for a group $h$. This means that $u_{h}$ has no effect on the conditional independence between $X_{j}$ and the other responses.
We consider the following convex problem:
where $\lambda\geq 0$ and $\lambda_{g}\geq 0$ are tuning parameters and $\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}=\sum_{h=1}^{q}\lVert\boldsymbol{\mathbf{\beta}}_{j,h}\rVert_{2}$ is the group lasso penalty. Together, the penalty on $\boldsymbol{\mathbf{\beta}}_{j}$ is the sparse-group lasso penalty developed in Simon et al. [2013]. The group lasso penalty is not applied $\boldsymbol{\mathbf{\beta}}_{j,0}$ nor to $\boldsymbol{\mathbf{\gamma}}_{j}$.
We considered the following two estimates of $\sigma^{2}_{\varepsilon_{j}}$,
$\hat{\sigma}^{2}_{1}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\hat{s}_{\beta_{j}}-\hat{s}_{\gamma_{j}}}\quad\text{and}\quad\hat{\sigma}^{2}_{2}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\hat{s}_{\beta_{j}}-1},$
where $\hat{s}_{\gamma_{j}}$ and $\hat{s}_{\beta_{j}}$ are the number of nonzero elements of $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ and $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$. Following the lasso literature [Reid et al., 2016, Yu and Bien, 2019], a straightforward estimate is given by $\hat{\sigma}^{2}_{1}$ as the value $\hat{s}_{\beta_{j}}+\hat{s}_{\gamma_{j}}$ roughly approximates the degrees of freedom of (15). In our experimentation, we have found $\hat{\sigma}^{2}_{2}$ to be slightly more consistent.
3.3 Implementation
We use block coordinate descent to solve (15) where the blocks are given by $\{\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j,0},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}\}$. Since (15) is convex and separable in these blocks, this approach is guaranteed to converge to the optimal solution [Tseng, 2001]. Our R implementation uses the sparsegl package to solve (15) [Liang et al., 2024]. We use $k$-fold cross-validation to select the tuning parameter pair $(\lambda,\lambda_{g})$. To do so, we set a parameter $\lambda_{0}>0$ and mixture parameter $\alpha_{s}\in[0,1]$ so that the penalty in (15) may be written as
Then for each $\alpha_{s}$, we run the method on a path of $\lambda_{0}$.

### v2

@@ 2.3 A convex formulation via natural parametrization
where $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})$ is the $p\times p$ diagonal matrix formed from $\boldsymbol{\mathbf{\Omega}}$. Under this parametrization, the nodewise regression model in (5) can be written as
@@
which yields a least squares criterion that is jointly convex in $\theta_{j}$ and $\boldsymbol{\mathbf{\Theta}}_{j,-j}$.
In contrast to (3), we model covariate dependence through $(\boldsymbol{\mathbf{\theta}},\boldsymbol{\mathbf{\Theta}})$ rather than $(\boldsymbol{\mathbf{\mu}},\boldsymbol{\mathbf{\Omega}})$ by specifying
$\boldsymbol{\mathbf{\theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h},$ (9)
where $\boldsymbol{\mathbf{B}}_{0}\in\mathbb{R}^{p\times p}$ represents the baseline graph structure and $\boldsymbol{\mathbf{B}}_{h}\in\mathbb{R}^{p\times p}$ captures the effect of covariate $u_{h}$ on the graph structure. Writing $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and letting $\beta_{jkh}=[\boldsymbol{\mathbf{B}}_{h}]_{jk}$ for $k\in[p]\setminus\{j\}$ and $h\in\{0\}\cup[q]$, this specification leads to the following nodewise regression model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}X_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}X_{k}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}),\quad\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$ (10)
for $j\in[p]$. That is, $X_{j}$ is regressed on the covariates $\boldsymbol{\mathbf{u}}$, all other responses $X_{k}$ for $k\neq j$, and the pairwise interactions between the covariates and the other responses. Remarkably, the least squares criterion associated with (10) is jointly convex in $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\beta_{jkh}$, enabling simultaneous estimation of all parameters and avoiding the nonconvexity inherent in the formulation of zhangli.
The parameterization (8) implicitly imposes that $\Theta_{jj}=-1$ for all $j\in[p]$. Under the specification (9), this is enforced by setting $\boldsymbol{\mathbf{B}}_{0}]_{jj}=-1$ and $\boldsymbol{\mathbf{B}}_{h}]_{jj}=0$ for all $h\in[q]$ and $j\in[p]$. Consequently, the residual variance $\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$, i.e., the conditional variance, does not depend on covariates, an assumption also adopted by zhangli. Moreover, $\boldsymbol{\mathbf{\Theta}}$ and $\boldsymbol{\mathbf{\Omega}}$ share the same sparsity structure, so that estimating the sparsity pattern of $\boldsymbol{\mathbf{\Theta}}$ is equivalent to that of $\boldsymbol{\mathbf{\Omega}}$. Accordingly, we focus on recovering the sparsity pattern of $\boldsymbol{\mathbf{\Theta}}$ in what follows.
Finally, we note that the proposed parametrization is closely related to, yet distinct from, existing convex formulations. The transformation in (8) is conceptually related to the natural parametrization in yu2019 for univariate-response regression and its multivariate extension in zhu_convex_2020. However, the scaling by $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}$ in (8) is crucial for ensuring that the nodewise regression model in (10) remains jointly convex in the presence of covariate adjustments. This feature distinguishes our approach from the reparametrizations in yu2019 and zhu_convex_2020, which do not accommodate covariate-dependent variance or covariance structures.
@@ 3 Estimation
For each $j\in[p]$, define the concatenated parameter vector $\boldsymbol{\mathbf{\beta}}_{j}=(\boldsymbol{\mathbf{\beta}}^{\top}_{j,0},\boldsymbol{\mathbf{\beta}}^{\top}_{j,1},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}^{\top})^{\top}\in\mathbb{R}^{(p-1)(q+1)}$, where the $h$-th block collects the parameters $\boldsymbol{\mathbf{\beta}}_{j,h}=(\beta_{j1h},\dotsc,\beta_{j,j-1,h},\beta_{j,j+1,h},\dotsc,\beta_{jph})^{\top}\in\mathbb{R}^{p-1}$. Under this notation, the nodewise regression model (10) can be written compactly as
@@
Figure 1 illustrates how each block $\boldsymbol{\mathbf{\beta}}_{j,h}^{\top}$ corresponds to the $j$-th row (and, by symmetry, the $j$-th column) of $\boldsymbol{\mathbf{\Theta}}$ in (9).
Suppose we collect $n$ independent observations $\{(\boldsymbol{\mathbf{x}}^{(i)},\boldsymbol{\mathbf{u}}^{(i)})\}_{i=1}^{n}$ of responses and covariates from the joint distribution in (1). Let $\boldsymbol{\mathbf{U}}=\left[\boldsymbol{\mathbf{u}}^{(1)},\dotsc,\boldsymbol{\mathbf{u}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times q}$ and $\boldsymbol{\mathbf{X}}=\left[\boldsymbol{\mathbf{x}}^{(1)},\dotsc,\boldsymbol{\mathbf{x}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times p}$ denote the covariate and response matrices, respectively. Let $\boldsymbol{\mathbf{u}}_{h}\in\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{x}}_{k}\in\mathbb{R}^{n}$ denote the $h$-th and $k$-th columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{X}}$. For $j\in[p]$, define the design matrix
$\boldsymbol{\mathbf{W}}_{-j}=\left[\boldsymbol{\mathbf{W}}_{-j,0},\ \boldsymbol{\mathbf{W}}_{-j,1},\ \dotsc,\ \boldsymbol{\mathbf{W}}_{-j,q}\right]\in\mathbb{R}^{n\times(p-1)(q+1)},$
where
$\begin{split}\boldsymbol{\mathbf{W}}_{-j,0}&=\left[\boldsymbol{\mathbf{x}}_{1},\dotsc,\boldsymbol{\mathbf{x}}_{j-1},\boldsymbol{\mathbf{x}}_{j+1},\dotsc,\boldsymbol{\mathbf{x}}_{p}\right]\in\mathbb{R}^{n\times(p-1)},\\ \boldsymbol{\mathbf{W}}_{-j,h}&=\left[\boldsymbol{\mathbf{x}}_{1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{j-1}\odot\boldsymbol{\mathbf{u}}_{h},\ \boldsymbol{\mathbf{x}}_{j+1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{p}\odot\boldsymbol{\mathbf{u}}_{h}\right]\in\mathbb{R}^{n\times(p-1)}\;\text{for $h\in[q]$},\end{split}$ (12)
and $\odot$ denotes the elementwise multiplication.
We consider estimating $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\boldsymbol{\mathbf{\beta}}_{j}$ in (11) by solving the following penalized least squares problem
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j}),$ (13)
where $g_{j}$ is a convex penalty function. In this paper, we focus on penalties that induce sparsity in $\boldsymbol{\mathbf{\beta}}_{j}$, promoting interpretable graph estimates. By construction, sparsity in $\boldsymbol{\mathbf{\beta}}_{j}$ corresponds to sparsity in the $j$-th row (and, by symmetry, the $j$-th column) of $\boldsymbol{\mathbf{\Theta}}$, and hence in $\boldsymbol{\mathbf{\Omega}}$ via (8).
@@
3.1 Sparse-group penalized estimation
Motivated by applications to eQTL studies, we consider a structured sparsity-inducing penalty for $g_{j}$ in (14). Our goal is to capture both elementwise and groupwise sparsity in the coefficients $\boldsymbol{\mathbf{\beta}}_{j}$. Elementwise sparsity within each group $\boldsymbol{\mathbf{\beta}}_{j,h}$ allows for the possibility that a covariate $u_{h}$ affects the conditional dependence between $X_{j}$ and only a subset of the remaining $\{X_{k}:k\neq j\}$. In contrast, groupwise sparsity, where $\boldsymbol{\mathbf{\beta}}_{j,h}=\boldsymbol{\mathbf{0}}$, indicates that the covariate $u_{h}$ has no effect on any conditional dependence structure involving $X_{j}$.
To this end, we consider the convex optimization problem
where $\lambda\geq 0$ and $\lambda_{g}\geq 0$ are tuning parameters, and $\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}=\sum_{h=1}^{q}\lVert\boldsymbol{\mathbf{\beta}}_{j,h}\rVert_{2}$ denotes the group lasso penalty [yuan2006model]. Together, the $\ell_{1}$ and group penalties on $\boldsymbol{\mathbf{\beta}}_{j}$ form a sparse-group lasso penalty [simon]. In contrast, only an $\ell_{1}$ penalty is applied to $\boldsymbol{\mathbf{\beta}}_{j,0}$ and $\boldsymbol{\mathbf{\gamma}}_{j}$, allowing for unrestricted elementwise sparsity in the baseline precision matrix $\boldsymbol{\mathbf{B}}_{0}$ and in the covariate effects on the mean vector.
To estimate the diagonal entries of $\boldsymbol{\mathbf{\Omega}}$, an estimate of the noise variance $\sigma^{2}_{\varepsilon_{j}}$ is required. While several estimators are available in the literature on error variance estimation in high-dimensional linear regression [reid, yu2019], we adopt
$\hat{\sigma}^{2}_{\varepsilon_{j}}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\widehat{\mathrm{df}}}$, (15)
where $\widehat{\mathrm{df}}$ denotes an estimate of the degrees of freedom of (14) as in li_highly_2018. Alternative approaches for estimating the degrees of freedom, or more generally the noise variance in (11), may also be employed.
3.2 Implementation details
We solve (14) using a blockwise coordinate descent algorithm, iteratively minimizing the objective with respect to each block in $\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j,0},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}$} while holding the others fixed. Each block update can be efficiently computed via a majorized proximal gradient method [simon, liang_sparsegl_2024]. Since (14) is convex and separable across blocks, the algorithm is guaranteed to converge to the optimal solution [Tseng].
The penalized nodewise regression problem (14) involves two tuning parameters: $\lambda$ controlling elementwise sparsity and $\lambda_{g}$ controlling groupwise sparsity. We select $(\lambda,\lambda_{g})$ via $k$-fold cross-validation. Equivalently, we reparametrize the penalty using a global sparsity penalty $\lambda_{0}>0$ and a mixing parameter $\alpha_{s}\in[0,1]$, so that the penalty in (14) becomes
For each fixed $\alpha_{s}$, we compute solution along a regularization path in $\lambda_{0}$.
As in standard nodewise regression approaches to GGMs, each row of the precision matrix is estimated separately, and the resulting estimate is not necessarily symmetric [meinshausen]. In our setting, each nodewise regression in (14) estimates one row of $\boldsymbol{\mathbf{\Theta}}$, and asymmetry arises due to both independent row estimation and the different row scaling factors in (8). To enforce symmetry, we apply a post hoc symmetrization step after performing all nodewise regressions. Specifically, with a solution $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$ to (14) and the corresponding error variance estimate in (15), set $\tilde{\boldsymbol{\mathbf{\beta}}}_{j}=-\hat{\boldsymbol{\mathbf{\beta}}}_{j}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ and construct symmetric matrices $\tilde{\boldsymbol{\mathbf{B}}}_{h}\in\mathbb{R}^{p\times p}$ for all $h=0,1,\dotsc,q$ with entries
@@ 3.2 Implementation details
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathds{1}\left[\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\right]+\tilde{\beta}_{kjh}\mathds{1}\left[\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\right$],
where the expression $\mathds{1}\left[P\right]$ is the indicator function of the argument $P$: it equals to $1$ if $P$ is true and 0 otherwise. This corresponds to the “and-rule” symmetrization, under which $\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero only if both $\tilde{\beta}_{jkh}$ and $\tilde{\beta}_{kjh}$ are nonzero. An alternative is the “or-rule,” defined by
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathds{1}{\left[\lvert\tilde{\beta}_{jkh}\rvert\geq\lvert\tilde{\beta}_{kjh}\rvert\right]}+\tilde{\beta}_{kjh}\mathds{1}{\left[\lvert\tilde{\beta}_{jkh}\rvert\leq\lvert\tilde{\beta}_{kjh}\rvert\right]}$,
which retains an edge if either coefficient is nonzero. We adopt the more conservative “and-rule”, noting that both approaches are asymptotically equivalent [meinshausen] and have been widely used in the GGM literature [cai2011constrained, liu2012tiger, Li_2021, tran2022completely, cai13, zhangli].
Finally, the mean vector and precision matrix are recovered via (8) as
@@
The full procedure is summarized in Algorithm 1 in Section S1.